### v1

# A Survey on Industrial Anomalies Synthesis

Abstract
Figure 1: (a) Trend of papers related to anomaly detection and anomaly synthesis from 2019 to 2024. Anomaly detection has shown steady growth in publications, while anomaly synthesis has gained increasing attention in recent years, particularly experiencing a significant surge in 2024. (b) Comparison of survey papers on anomaly detection and anomaly synthesis. While numerous surveys exist on anomaly detection, anomaly synthesis remains an emerging field with no dedicated surveys to date. Our work is the first to fill this gap. Data sources are from https://github.com/M-3LAB/awesome-industrial-anomaly-detection
This paper comprehensively reviews anomaly synthesis methodologies. Existing surveys focus on limited techniques, missing an overall field view and understanding method interconnections. In contrast, our study offers a unified review, covering about 40 representative methods across Hand-crafted, Distribution-hypothesis-based, Generative models (GM)-based, and Vision-language models (VLM)-based synthesis. We introduce the first industrial anomaly synthesis (IAS) taxonomy. Prior works lack formal classification or use simplistic taxonomies, hampering structured comparisons and trend identification. Our taxonomy provides a fine-grained framework reflecting methodological progress and practical implications, grounding future research. Furthermore, we explore cross-modality synthesis and large-scale VLM. Previous surveys overlooked multimodal data and VLM in anomaly synthesis, limiting insights into their advantages. Our survey analyzes their integration, benefits, challenges, and prospects, offering a roadmap to boost IAS with multimodal learning. More resources are available at https://github.com/M-3LAB/awesome-anomaly-synthesis.
1 Introduction
Image anomaly detection plays a pivotal role in manufacturing, primarily focused on identifying the anomalies of products, thereby ensuring product quality. In practical usage, an image anomaly detection system typically requires substantial quantities of high-quality annotated abnormal samples for training. However, the acquisition of high-quality annotated abnormal samples incurs prohibitively high costs. The difficulties are summarized as follows: ❶ The low defective rate presents a fundamental challenge. In large-scale manufacturing scenarios, the proportion of qualified products far exceeds that of abnormal ones. ❷ The necessity for specialized equipment substantially escalates costs. Many industrial anomalies, such as microscopic cracks, fine scratches, or concealed contaminants, require detection via high-precision instruments (e.g. high-magnification microscopes, and X-ray inspection systems). These high-precision machines advance technical infrastructure investment. ❸ Abnormal sample labeling requires domain expertise and meticulous analysis, resulting in time-consuming procedures. Accurate annotation requires skilled professionals, whose limited availability increases labor expenses. Additionally, certain anomalies demand multi-modal validation using advanced imaging techniques, further amplifying the time and resource burden.
To address these limitations, various image anomalies synthesis algorithms have been developed to generate abnormal samples, aiming to mitigate the scarcity of real abnormal samples and enhance downstream detection performances. Nevertheless, the existing anomalies synthesis algorithm cannot fulfill the requirements of industrial manufacturing. The challenges are listed as follows:
  • 1.
    Limited Sampling of Anomalies Distribution The finite number of abnormal samples for specific anomaly types constitutes sparse sampling from the underlying anomaly distribution, failing to capture full variability. This constraint forces the algorithm to adopt increased complexity to satisfy diversity requirements.
  • 2.
    Realistic Anomalies Synthesis Practical anomalies are highly complex (e.g., cracks, scratches, contaminants, punctures) and exhibit large distribution shifts compared to background textures. The diverse occurrence rates of different anomaly types add significant complexity to the modeling process. It can lead to synthetic anomalies that fail to accurately capture the realism of real-world anomalies, resulting in missing or unrealistic features.
  • 3.
    Underutilization of Multimodal Information While multimodal cues (e.g. text prompts for anomalies synthesis) are increasingly available, effectively integrating such information to synthesize realistic anomalies patterns remains an open challenge.
Consequently, systematic analysis and comprehensive review of current industrial anomaly synthesis (IAS) have become imperative. As shown in Figure 1, the number of published papers on anomaly detection has surged significantly in recent years, highlighting the growing industrial and academic focus on this field. Correspondingly, research on anomaly synthesis has also seen a noticeable increase, reflecting the rising recognition of its importance in mitigating the scarcity of real abnormal samples. Despite this upward trend, existing anomaly synthesis methods still face fundamental challenges that hinder their practical adoption in industrial manufacturing. By critically examining these challenges, we can not only identify the strengths and limitations of current research but also establish clear guidelines for future algorithm development. Such an analysis will provide foundational insights for advancing anomaly synthesis techniques and optimizing their applicability in real-world scenarios.
Comparison to Existing Surveys: Therefore, we propose a comprehensive survey that not only bridges the gaps in existing work but also establishes a novel framework for systematic analysis in IAS. Although several surveys have briefly discussed IAS, they exhibit specific limitations shown in Table. 1, which can also be categorized as follows:
(i) Limited Scope on Synthesis Methods. Existing surveys often limit their scope to specific synthesis methods, failing to provide a holistic perspective on the field. For instance,  Chen et al. emphasize data augmentation strategies but overlook the potential of more advanced synthesis methods in generating more discriminative defect representations. Similarly,  Xia et al. focus on the performance of generative adversarial networks (GAN)-based models in anomaly detction and synthesis, but fail to explore the potential of other emerging generative paradigms. This fragmented perspective restricts the understanding of how diverse methods can complement each other in addressing practical challenges.
(ii) Lack of a Dedicated Taxonomy: Existing anomaly-related surveys primarily focus on anomaly detection, often treating anomaly synthesis as a minor aspect without a structured classification. For instance,  Liu et al. broadly classify anomaly detection methods into supervised and unsupervised categories, which fails to capture the methodological diversity of anomaly synthesis. To the best of our knowledge, our survey is the first dedicated work that systematically categorizes IAS methods. By addressing this gap, we introduce a structured taxonomy tailored specifically for anomaly synthesis, providing a more comprehensive foundation for understanding and comparing different synthesis approaches.
(iii) Insufficient Emphasis on Cross-Modality Synthesis. Although existing research acknowledges the importance of leveraging multiple modalities, many review papers have not thoroughly explored how emerging technologies, such as large-scale vision-language models (VLM), are transforming the field of anomaly synthesis.  Cao et al. proposed the use of VLM for cross-modal anomaly synthesis. However, a systematic analysis of their applicability and current limitations remains urgently needed, the lack of which has hindered the exploration of how advanced methods align with the practical demands of industry.
Table 1: Comparison of previous survies and our survey.
 Liu et al.  Chen et al.  Xia et al.  Cao et al. Our
Perspective Detection Detection Detection Detection Synthesis
Number of IAS Categories 2 0 4 4 10
Multimodal Interaction
Figure 2: A Taxonomy of Image Anomalies Synthesis(IAS)
Contributions: The main contributions of this survey paper can be summarized as follows:
  • 1.
    Our survey provides a unified and systematic review of anomaly synthesis, covering nearly 40 representative methods across different paradigms. By exploring Hand-crafted synthesis, Distribution hypothesis-based synthesis, Generative models (GM)-based synthesis, and Vision language models (VLM)-based synthesis, we provide a comprehensive overview that captures the full scope of techniques available in the field.
  • 2.
    We present the first dedicated taxonomy for IAS, offering a structured and fine-grained classification framework that reflects methodological advancements and practical implications, thus serving as a foundation for future research and innovation in IAS.
  • 3.
    This work delves into the integration of multimodal cues and large-scale VLM in anomaly synthesis, addressing their potential benefits, challenges, and future opportunities. Our exploration provides a comprehensive framework for leveraging multimodal learning, enhancing the effectiveness of anomaly synthesis in industrial applications.
Organization: The remainder of this paper is structured as follows: Section 2 presents a systematic taxonomy of IAS4, introducing key paradigms and their methodological distinctions. Sections 3 to 6 provide an in-depth analysis of the four primary categories of IAS, examining their theoretical foundations, implementation strategies, and practical applications. Specifically, these sections cover Hand-crafted synthesis, Distribution hypothesis-based synthesis, Generative models (GM)-based synthesis, and recently emerging Vision language models (VLM)-based synthesis. Finally, Sections 7 and 8 consolidate key insights from the survey, critically assess the limitations of current methodologies, and outline promising future research directions to advance the field of IAS.
2 Taxonomy of IAS
Taxonomy of Industrial Anomaly Synthesis: In Figure 2, we present a detailed taxonomy of IAS. We categorize IAS into four main paradigms: Hand-crafted synthesis, Distribution hypothesis-based synthesis, Generative models (GM)-based synthesis, and Vision language models (VLM)-based synthesis. Each paradigm has distinct characteristics in terms of the anomaly synthesis approach and applicable scenarios. Additionally, Figure 3 illustrates the internal structures and implementation details of these four paradigms, providing a comprehensive view of the anomaly synthesis process within the different approaches.
Definition of Hand-Crafted Synthesis: Hand-crafted synthesis relies on manually designed rules to simulate anomalies, typically training-free and suitable for controlled environments where high realism and defect diversity are not critical. Self-contained synthesis manipulates the original image through operations like cropping or rearrangement to simulate texture misalignment or color variation, synthesizing abnormal regions. External-dependent synthesis employs external data (e.g., texture libraries) to synthesize anomalies independently of the original image, ensuring abnormal parts are not confined by its content. Inpainting-based syhthesis remove local areas through masking techniques, disrupting structural continuity by adding noise or black patches to generate anomalies.
Definition of Distribution Hypothesis-based Synthesis: Distribution hypothesis-based synthesis relies on statistical modeling of normal data distributions and synthesize anomalies through controlled perturbations, typically by adjusting the feature space of normal samples. Prior-dependent synthesis uses the pre-defined geometric assumptions (e.g., manifold or hypersphere structures) to define normal data distributions in feature space, applying controlled deviations to ensure the synthesized feature-level anomalies lie at the boundary or outside the normal distribution. In contrast, Data-driven synthesis leverages intrinsic statistical properties of data by extracting features in the latent space and synthesizing anomalies through perturbations or adaptive strategies, enhancing the diversity and realism of the generated samples.
Definition of GM-based Synthesis: Recent advancements in deep GM, such as GANs and diffusion methods, enable realistic anomaly synthesis. GM-based synthesis is categorized into Full-image synthesis, Full-image translation, and Local anomalies synthesis. Full-image synthesis learns abnormal data distributions and constructs the mapping from random noise to abnormal samples. Full-image translation uses domain translation techniques to map normal images to abnormal ones, injecting anomalies while preserving the global structure. Local anomalies synthesis replaces specific regions of normal images with learned local anomalies and ensures smooth transitions between abnormal areas and the background.
Definition of VLM-based Synthesis: Leveraging large-scale pre-trained VLMs with billions of parameters, VLM-based synthesis exploits extensive pre-trained knowledge and integrated multimodal cues to synthesize high-quality anomalies. Single-stage synthesis directly produces realistic, context-aware, and detailed abnormal samples. It applies prompt engineering or lightweight fine-tuning for computationally efficient anomaly synthesis. In contrast, multi-stage synthesis employs a complete pipeline that refines both global and local features—integrating synthetic abnormal data with mask synthesis or other multi-process optimization techniques—to enhance realism, diversity, and downstream task alignment.
3 Hand-crafted Synthesis
Figure 3: Illustration of different IAS methods. It categorizes IAS methods into four main approaches: (a) Hand-crafted synthesis, which relies on pre-defined textures such as external-dependent synthesis, self-contained synthesis, and inpainting-based synthesis. (b) Distribution hypothesis-based synthesis, which synthesizes anomalies by perturbing learned feature distributions of normal data through prior-dependent or data-driven synthesis. (c) Generative models (GM)-based synthesis, which employs generative models to perform full-image generation, full-image translation, or local anomaly synthesis. (d) Vision-language models (VLM)-based synthesis, which integrates textual features into a single-stage or multi-stage synthesis process.
Self-contained synthesis is a vital subset of hand-crafted synthesis. It operates by directly manipulating image regions, such as cropping, rearranging, or perturbing portions of the image, to synthesize new anomalies derived entirely from the original image. For instance,  Li et al. and  Schlüter et al. propose randomly cropping rectangular patches and pasting them back into normal images. Additionally,  Lyu et al. use Bézier curves to define the shape of anomalies, and CutPaste-based augmentation is employed to synthesize abnormal regions with the guidance of a saliency model. Moreover,  Pei et al. extract patches from the same object, apply diverse augmentations (e.g., geometric transformations, color distortions), and paste them back onto the object using a mask-guided strategy, enhancing downstream anomaly detection performance.
Strengths and Weaknesses. Self-contained synthesis offers a straightforward and cost-efficient approach by generating anomalies solely from the content of the original images. By directly manipulating regions within an image—such as cropping or rearranging—it maintains the original structural and textural context, thereby enhancing the model’s sensitivity to nuanced anomalies. However, since the anomalies derive exclusively from existing image features, the range and complexity of anomalies that it can simulate are inherently limited, which can result in a reduced ability to capture the full spectrum of real-world anomaly variations, potentially affecting the performance of downstream models when faced with more complex or diverse anomalies.
External-dependent synthesis is an approach that synthesizes anomalies by incorporating textures from external data, such as texture libraries. Unlike self-contained methods, it overlays or blends external areas onto an otherwise anomaly-free image. For example,  Zavrtanik et al. and  Zhang et al. use Perlin noise to generate binary anomaly masks, enabling the combination of clean backgrounds with external textures to produce abnormal samples. Similarly,  Yang et al. refine the process by restricting the location of anomalies to the foreground, which further bridges the gap between synthetic and real anomalies.
Strengths and Weaknesses. External-dependent synthesis can produce a wider variety of anomaly patterns by leveraging external data. It allows the generation of anomalies with distinct distributions that may be challenging to replicate using only intrinsic image content, thus potentially improving the robustness of detection models. However, integrating external textures seamlessly with the original content can be difficult, often resulting in inconsistencies or visual artifacts, particularly when high-fidelity anomaly synthesis is crucial.
Inpainting-based synthesis is a specialized category within hand-crafted synthesis that synthesizes anomalies by selectively masking regions of the original image, disrupting its visual continuity [Pirnay and Chai]. Unlike methods that paste self-contained content or incorporate external textures, inpainting-based synthesis focuses on introducing missing or occluded areas, often by blacking out regions or filling them with noise. This transformation from a complete to an incomplete image is particularly effective in downstream reconstruction-based anomaly detection. For instance,  Nakanishi et al.,  Zavrtanik et al., and Li et al. synthesize abnormal samples by masking random regions and then reconstructing them using background information. In contrast,  Luo et al. develop an adaptive mask generator that selectively conceals abnormal regions while preserving the surrounding normal background, resulting in a more structured and context-aware synthesis. Additionally, Cao et al. and Tien et al. introduce random noise within the masked regions.
Strengths and Weaknesses. Inpainting-based synthesis is a straightforward approach that involves replacing some areas of an image with black patches or noise to simulate anomalies. It effectively evaluates the performance of reconstruction-based detection models. However, the generated anomalies tend to be overly simplistic and may lack the complexity and subtle nuances characteristic of real-world anomalies.
4 Distribution hypothesis-based Synthesis
Prior-dependent synthesis is a core approach in Distribution hypothesis-based syhthesis that leverages a geometric prior. It assumes that the latent space of normal data maps within a defined region, typically a manifold or hypersphere, while anomalies fall outside this boundary. By modeling the normal feature distribution and applying controlled perturbations, synthesized anomalies are at or beyond these borders, effectively simulating real anomaly features and improving the preformance of downstream task. For instance, Chen et al. constrain normal features within a compact space and guide synthesis via gradient ascent and truncated projection for enhanced diversity. Similarly, Chen et al. generate anomalies radially to refine decision boundaries, while Shin et al. utilize the manifold hypothesis and Gaussian annulus to perturb features and compute restoration errors for localization. In addition, Naud and Lavin embed normal data in non-Euclidean spaces and perturb along geodesic paths to synthesize anomalies.
Strengths and Weaknesses. Prior-dependent synthesis leverages latent assumptions from normal samples to enable controlled perturbations, enhancing downstream tasks like classification or detection. Modeling normal data within a defined latent space (e.g., a manifold or hypersphere) facilitates these perturbations. However, as most methods synthesize anomalies at the feature level, they lack spatial details for precise anomaly segmentation. Moreover, reliance on a predefined latent distribution can limit the ability to capture complex spatial anomalies, reducing broader applicability.
Data-driven synthesis is an effective approach within Distribution Hypothesis-Based Models that synthesizes anomalies by directly manipulating latent representations of normal data. Instead of relying on explicit prior assumptions, this method extracts latent features using models like autoencoders and then introduces controlled perturbations, such as Gaussian noise or other data-adaptive constraints, to synthesize anomalies. This flexible strategy leverages the intrinsic statistical properties of the data, producing anomalies that closely mirror distributions of abnormal features. Recent advancements explore diverse approaches to enhance the flexibility and realism of data-driven synthesis.  Liu et al. and  You et al. introduce SimpleNet and UniAD, which synthesize anomalies directly at the feature level by adding noise to extracted features to simulate abnormal features. The perturbed and original features in SimpleNet are then evaluated by a discriminator to ensure that the synthetic anomalies are distinct yet plausible. Building upon this,  Rolih et al. propose SuperSimpleNet, which confines the noise to specific regions and employs a new segmentation head, offering a more targeted anomaly synthesis strategy. Additionally,  Zavrtanik et al. develop DSR, a dual subspace re-projection network that trains a codebook and replaces the contents of masked regions in normal features by sampling from this codebook to synthesize abnormal features.
Strengths and Weaknesses. Data-driven synthesis methods avoid explicit distribution assumptions by directly learning latent representations of normal samples. By mapping normal data in latent space via neural networks and applying perturbations, it synthesizes diverse anomaly features resembling those in the real world. However, its effectiveness relies on the quality of the latent space, since poorly trained models may produce unrealistic anomalies. Additionally, without accurately estimating the normal distribution, the generated anomalies may lead to suboptimal decision boundaries.
5 GM-based Synthesis
Full-image synthesis is a fundamental approach within GMs-Based Models, extensively employed for directly synthesizing industrial anomalies. It utilizes GM such as GAN and diffusion models to construct a unique mapping that transforms Gaussian noise into abnormal samples, effectively approximating the distribution of real anomalies. By training distinct models for different types of anomalies, it becomes possible to generate a wide variety of high-quality data that closely to actual anomalies.  Liu et al. establish multistage GANs that decouple texture generation and background-anomaly fusion, effectively modeling both abnormal areas and contextual coherence. To address data scarcity,  Du et al. develop Con-GAN, featuring shared data augmentation and hypersphere-based loss, it simultaneously prevent overfitting and enhance anomaly diversity.  Duan et al. also propose a two-stage GAN, namely DFMGAN, which consists of two processes: initial learning on normal samples and subsequent fine-tuning using abnormal samples. For complex anomaly synthesis,  Yang et al. innovatively combine large receptive fields for global structural modeling with patch-level refinement mechanisms, achieving synergistic integration of macro-context and micro-details.
Strengths and Weaknesses. While full-image synthesis effectively enhances abnormal sample quality, its performance is highly dependent on the availability and diversity of training data. Limited abnormal samples often lead to artifacts or unrealistic defects, restricting practical applicability. Additionally, as these models generate entire images rather than directly modifying normal samples, they struggle to preserve fine-grained structural details and contextual consistency.
Full-image translation is a key approach within GM-Based Models, it synthesizes anomalies by transforming normal images via domain translation techniques like CycleGAN [Zhu et al.] and Pix2PixGAN [Isola et al.]. Full-image translation learns mappings between normal and anomaly domains, introducing targeted modifications (e.g. scratches, stains) while preserving structural integrity. Nowadays, it is a highly effective strategy for IAS. Recent advancements significantly improve full-image translation by enhancing the fidelity and controllability of synthesized anomalies.  Wen et al. extend CycleGAN within their anomaly detection pipeline, achieving superior performance.  Niu et al. propose SDGAN, a CycleGAN-based framework incorporating additional discriminators to refine the translation process. By focusing on modeling the distribution of abnormal samples, SDGAN enhances the realism of synthesized anomalies. Furthermore,  Zhang et al. introduce Defect-GAN, which explicitly models both the defacement and restoration processes rather than synthesizing anomalies arbitrarily. Additionally, it leverages spatial distribution maps to preserve the appearance of normal backgrounds, ensuring structural consistency in generated images.
Strengths and Weaknesses. Full-image translation excels at generating realistic, context-aware anomalies with minimal abnormal samples and numerous normal samples. Its strength lies in seamlessly integrating anomaly into normal contexts. However, its control over anomaly types and spatial distribution is limited. The inherent property of Full-image translation makes it challenging to precisely dictate where and how anomalies appear, potentially reducing their applicability in scenarios requiring fine-grained anomaly localization.
Local anomaly synthesis is a targeted approach within GMs-Based Models, designed to generate localized anomalies by selectively modifying specific regions of an image. Local anomaly synthesis typically involves removing parts of the image through annotations and employing generative models to synthesize anomaly-like textures within the masked regions. By focusing on localized transformations, it enables the creation of controlled and diverse anomaly patterns, such as scratches, holes, or surface irregularities, which closely resemble real-world anomalies. Recent advancements enhance local anomaly synthesis by aligning the mask with synthesized abnormal regions and improving spatial consistency between synthetic defects and their surrounding contexts.  Zhang et al. enhance the synthesis of realistic anomalies by amplifying their divergence from normal patterns. Their proposed RealNet perturbs the variance of denoising diffusion during reverse diffusion to synthesize a global anomaly map.  Wei et al. construct pseudo-normal backgrounds in abnormal regions to emphasize the distribution of anomalies.  Wei et al. develop DCDGAN, which trains a model on anomaly-only textures and blends them with different backgrounds, achieving multi-class and diversified anomaly synthesis.
Strengths and Weaknesses. Local anomaly synthesis excels at synthesizing controlled abnormal samples, making it suitable for targeted anomaly synthesis. Restricting anomalies to specific regions addresses the impact of synthetic backgrounds on downstream tasks, which is a common issue in GM-based approaches. However, it struggles to maintain texture-background coherence. Additionally, it depends on mask annotations to define abnormal regions, which makes acquiring high-quality annotations costly and labor-intensive, limiting scalability in large-scale industrial applications.
6 VLM-based Synthesis
Single-stage synthesis leverages pre-trained VLMs with billions of parameters, leveraging their abundant feature representations and multimodal information. It synthesizes realistic anomalies and produces context-aware, highly detailed abnormal samples with minimal or no adjustments. By capitalizing on abundant cues from various modalities, single-stage synthesis balances computational efficiency and domain specificity, making it an attractive solution for industrial applications.  Sun et al. introduce a training-free anomaly synthesis method based on stable diffusion that maximizes alignment between abnormal regions and anomaly descriptions.  Hu et al. generate embeddings from diverse modalities and adjust them dynamically during the diffusion process. This strategy enables the creation of anomalies enriched with multimodal information, resulting in higher fidelity and diversity. Additionally,  Jiang et al. propose CAGEN, which employs mask annotations and text prompts to fine-tune ControlNet for precise control over the location and type of synthesized anomalies. Furthermore,  He et al. model cross-modal semantic features to allow fine-grained control over the characteristics of generated anomalies.
Strengths and Weaknesses. Single-stage synthesis harnesses diverse multimodal cues to synthesize realistic abnormal textures. By leveraging both visual and textual modalities, it effectively captures detailed information, enabling more precise anomaly synthesis while allowing fine-grained control through prompt modifications. However, a significant limitation is that the synthesized anomaly regions often do not precisely align with mask annotations, which can adversely impact the accuracy of downstream anomaly segmentation tasks. This misalignment also leads to inconsistencies between the abnormal data and annotations, reducing reliability in industrial applications.
Multi-stage synthesis is an advanced VLM technique that provides a comprehensive pipeline for anomaly synthesis. It integrates multiple processes, including anomaly synthesis and precise mask production. By iteratively refining global context and local abnormal regions, it enhances the realism, diversity, and contextual consistency of synthesized anomalies while meeting the requirements of downstream tasks better. This complete pipeline strengthens the usability and reliability of industrial systems. Recent studies contribute notably to this field.  Jin et al. propose a dual-branch VLM where one branch captures global context and the other targets abnormal regions, and it leverages a pre-trained segmentation for high-quality mask generation.  Shi et al. integrate textual guidance into global information and local synthesis, introducing a novel mask algorithm that improves small anomaly synthesis and enables precise control over intensity and direction.  Dai et al. introduce SeaS, a few-shot method that binds anomaly attributes to specific tokens using unbalanced anomaly prompts while aligning normal image features to preserve authenticity, it can also produce accurate mask annotations. Lastly,  Hu et al. develop AnomalyDiffusion, incorporating textual and positional information for precise anomaly localization and employing textual inversion to learn mask embeddings.
Strengths and Weaknesses. Multi-stage synthesis integrates sequential training and precise anomaly mask generation to refine anomaly synthesis. It can capture global context and local details, enhancing the fidelity and diversity of the synthesized anomalies. The advanced mask generation enables anomaly synthesis at designated locations, further aligning with industrial needs. However, its complexity and high computational cost limit scalability in resource-constrained settings. Despite this, multi-stage synthesis remains essential for applications demanding realistic, context-aware anomalies in specific industrial scenarios.
7 Future dirction
Enhance Anomaly Diversity. The limited coverage of anomaly distributions remains a critical bottleneck in IAS. Existing methods often rely on a limited set of real anomalies, leading to overfitting and insufficient diversity in synthesized anomalies. A potential approach is adaptive anomaly synthesis techniques leveraging self-supervised and active learning algorithms. For instance, uncertainty-aware models can guide synthesis by identifying underrepresented abnormal regions and dynamically adjusting it. Additionally, a coarse-to-fine approach, first generating structures then refining details, can improve diversity. Furthermore, constructing larger, more diverse datasets remains crucial for better generalization.
Controllable Synthesis of Anomaly Attributes. Industrial image anomalies exhibit substantial complexity, including variations in shape, texture, and distribution. Existing approaches often struggle to handle these variations, leading to the generation of unrealistic or overly simplified anomalies. To improve this, future IAS researches require to focus on cross-class consistency modeling, leveraging advanced network architectures, such as large-scale models, to learn shared and distinct characteristics across different anomaly types. Additionally, controllable synthesis techniques should be employed to allow precise adjustments of anomaly attributes like type and location. For instance, combining local anomaly synthesis with advanced segmentation models facilitates fine-tuned modifications of anomaly attributes, ensuring seamless integration with background and improve the realism of synthesized samples.
Promote Multimodal Anomaly Synthesis. Despite the increasing availability of complementary modalities such as textual descriptions, infrared imagery, and X-ray scans, their integration into IAS remains largely unexplored. Future advancements should focus on developing cross-modal alignment strategies that leverage VLM and multimodal transformers to establish semantic correlations between different modalities, enabling richer anomaly synthesis. Furthermore, incorporating contrastive learning techniques can improve cross-modal feature learning, ensuring better correspondence between generated anomalies and their real-world counterparts. Multi-source data fusion techniques can also be explored to enhance the robustness of synthesized anomalies, leveraging complementary strengths of different modalities. And integrating reinforcement learning into multimodal synthesis pipelines could enable adaptive anomaly generation, dynamically optimizing synthesis strategies based on anomaly characteristics and industrial requirements.
8 Conclusion
In this survey, we explored recent advancements in Industrial Image Anomaly Synthesis (IAS) by first identifying three key challenges: ❶ limited sampling in anomalies distribution, ❷ difficulty in synthesizing realistic anomalies, and ❸ underutilization of multimodal information. To address them, we provided essential background knowledge on IAS and introduced its core methodologies. We then conducted a comprehensive review of existing IAS approaches, categorizing them into Hand-crafted, Distribution hypothesis-based, Generative models-based, and Vision language models-based synthesis. Finally, we discussed promising future research directions, such as enhancing anomaly diversity, achieving a controllable synthesis of anomaly attributes, and fully leveraging multimodal information. We believe that tackling these challenges will significantly improve the effectiveness of IAS.

### v2

@@
 Existing surveys often limit their scope to specific synthesis methods, failing to provide a holistic perspective on the field.
-For instance,  Chen et al.
-emphasize data augmentation strategies but overlook the potential of more advanced synthesis methods in generating more discriminative defect representations.
-Similarly,  Xia et al.
-focus on the performance of generative adversarial networks (GAN)-based models in anomaly detction and synthesis, but fail to explore the potential of other emerging generative paradigms.
+For instance,  chen2021surface emphasize data augmentation strategies but overlook the potential of more advanced synthesis methods in generating more discriminative defect representations.
+Similarly,  xia2022gan focus on the performance of generative adversarial networks (GAN)-based models in anomaly detction and synthesis, but fail to explore the potential of other emerging generative paradigms.
 This fragmented perspective restricts the understanding of how diverse methods can complement each other in addressing practical challenges.
 (ii) Lack of a Dedicated Taxonomy: Existing anomaly-related surveys primarily focus on anomaly detection, often treating anomaly synthesis as a minor aspect without a structured classification.
-For instance,  Liu et al.
-broadly classify anomaly detection methods into supervised and unsupervised categories, which fails to capture the methodological diversity of anomaly synthesis.
+For instance,  liu2024deep broadly classify anomaly detection methods into supervised and unsupervised categories, which fails to capture the methodological diversity of anomaly synthesis.
 To the best of our knowledge, our survey is the first dedicated work that systematically categorizes IAS methods.
@@
 Although existing research acknowledges the importance of leveraging multiple modalities, many review papers have not thoroughly explored how emerging technologies, such as large-scale vision-language models (VLM), are transforming the field of anomaly synthesis.
-Cao et al.
-proposed the use of VLM for cross-modal anomaly synthesis.
+cao2024survey proposed the use of VLM for cross-modal anomaly synthesis.
 However, a systematic analysis of their applicability and current limitations remains urgently needed, the lack of which has hindered the exploration of how advanced methods align with the practical demands of industry.
 Table 1: Comparison of previous survies and our survey.
- Liu et al.  Chen et al.  Xia et al.  Cao et al. Our
+ liu2024deep  chen2021surface  xia2022gan  cao2024survey Our
 Perspective Detection Detection Detection Detection Synthesis
@@
 Figure 2: A Taxonomy of Image Anomalies Synthesis(IAS)
-Contributions: The main contributions of this survey paper can be summarized as follows:
   • 1.
@@
 It operates by directly manipulating image regions, such as cropping, rearranging, or perturbing portions of the image, to synthesize new anomalies derived entirely from the original image.
-For instance,  Li et al.
-and  Schlüter et al.
-propose randomly cropping rectangular patches and pasting them back into normal images.
-Additionally,  Lyu et al.
-use Bézier curves to define the shape of anomalies, and CutPaste-based augmentation is employed to synthesize abnormal regions with the guidance of a saliency model.
-Moreover,  Pei et al.
-extract patches from the same object, apply diverse augmentations (e.g., geometric transformations, color distortions), and paste them back onto the object using a mask-guided strategy, enhancing downstream anomaly detection performance.
+For instance,  li2021cutpaste and  schluter2022natural propose randomly cropping rectangular patches and pasting them back into normal images.
+Additionally,  lyu2024reb use Bézier curves to define the shape of anomalies, and CutPaste-based augmentation is employed to synthesize abnormal regions with the guidance of a saliency model.
+Moreover,  pei2023self extract patches from the same object, apply diverse augmentations (e.g., geometric transformations, color distortions), and paste them back onto the object using a mask-guided strategy, enhancing downstream anomaly detection performance.
 Strengths and Weaknesses.
@@
 Unlike self-contained methods, it overlays or blends external areas onto an otherwise anomaly-free image.
-For example,  Zavrtanik et al.
-and  Zhang et al.
-use Perlin noise to generate binary anomaly masks, enabling the combination of clean backgrounds with external textures to produce abnormal samples.
-Similarly,  Yang et al.
-refine the process by restricting the location of anomalies to the foreground, which further bridges the gap between synthetic and real anomalies.
+For example,  zavrtanik2021draem and  zhang2023destseg use Perlin noise to generate binary anomaly masks, enabling the combination of clean backgrounds with external textures to produce abnormal samples.
+Similarly,  yang2023memseg refine the process by restricting the location of anomalies to the foreground, which further bridges the gap between synthetic and real anomalies.
 Strengths and Weaknesses.
@@
 However, integrating external textures seamlessly with the original content can be difficult, often resulting in inconsistencies or visual artifacts, particularly when high-fidelity anomaly synthesis is crucial.
-Inpainting-based synthesis is a specialized category within hand-crafted synthesis that synthesizes anomalies by selectively masking regions of the original image, disrupting its visual continuity [Pirnay and Chai].
+Inpainting-based synthesis is a specialized category within hand-crafted synthesis that synthesizes anomalies by selectively masking regions of the original image, disrupting its visual continuity [pirnay2022inpainting].
 Unlike methods that paste self-contained content or incorporate external textures, inpainting-based synthesis focuses on introducing missing or occluded areas, often by blacking out regions or filling them with noise.
 This transformation from a complete to an incomplete image is particularly effective in downstream reconstruction-based anomaly detection.
-For instance,  Nakanishi et al.,  Zavrtanik et al., and Li et al.
-synthesize abnormal samples by masking random regions and then reconstructing them using background information.
-In contrast,  Luo et al.
-develop an adaptive mask generator that selectively conceals abnormal regions while preserving the surrounding normal background, resulting in a more structured and context-aware synthesis.
-Additionally, Cao et al.
-and Tien et al.
-introduce random noise within the masked regions.
+For instance,  nakanishi2021iterative,  zavrtanik2021reconstruction, and li2020superpixel synthesize abnormal samples by masking random regions and then reconstructing them using background information.
+In contrast,  luo2024ami develop an adaptive mask generator that selectively conceals abnormal regions while preserving the surrounding normal background, resulting in a more structured and context-aware synthesis.
+Additionally, cao2023collaborative and tien2023revisiting introduce random noise within the masked regions.
 Strengths and Weaknesses.
@@
 By modeling the normal feature distribution and applying controlled perturbations, synthesized anomalies are at or beyond these borders, effectively simulating real anomaly features and improving the preformance of downstream task.
-For instance, Chen et al.
-constrain normal features within a compact space and guide synthesis via gradient ascent and truncated projection for enhanced diversity.
-Similarly, Chen et al.
-generate anomalies radially to refine decision boundaries, while Shin et al.
-utilize the manifold hypothesis and Gaussian annulus to perturb features and compute restoration errors for localization.
-In addition, Naud and Lavin embed normal data in non-Euclidean spaces and perturb along geodesic paths to synthesize anomalies.
+For instance, chen2025unified constrain normal features within a compact space and guide synthesis via gradient ascent and truncated projection for enhanced diversity.
+Similarly, chen2024progressive generate anomalies radially to refine decision boundaries, while shin2023anomaly utilize the manifold hypothesis and Gaussian annulus to perturb features and compute restoration errors for localization.
+In addition, naud2020manifolds embed normal data in non-Euclidean spaces and perturb along geodesic paths to synthesize anomalies.
 Strengths and Weaknesses.
@@
 Recent advancements explore diverse approaches to enhance the flexibility and realism of data-driven synthesis.
-Liu et al.
-and  You et al.
-introduce SimpleNet and UniAD, which synthesize anomalies directly at the feature level by adding noise to extracted features to simulate abnormal features.
+liu2023simplenet and  you2022unified introduce SimpleNet and UniAD, which synthesize anomalies directly at the feature level by adding noise to extracted features to simulate abnormal features.
 The perturbed and original features in SimpleNet are then evaluated by a discriminator to ensure that the synthetic anomalies are distinct yet plausible.
-Building upon this,  Rolih et al.
-propose SuperSimpleNet, which confines the noise to specific regions and employs a new segmentation head, offering a more targeted anomaly synthesis strategy.
-Additionally,  Zavrtanik et al.
-develop DSR, a dual subspace re-projection network that trains a codebook and replaces the contents of masked regions in normal features by sampling from this codebook to synthesize abnormal features.
+Building upon this,  rolih2025supersimplenet propose SuperSimpleNet, which confines the noise to specific regions and employs a new segmentation head, offering a more targeted anomaly synthesis strategy.
+Additionally,  zavrtanik2022dsr develop DSR, a dual subspace re-projection network that trains a codebook and replaces the contents of masked regions in normal features by sampling from this codebook to synthesize abnormal features.
 Strengths and Weaknesses.
@@
 By training distinct models for different types of anomalies, it becomes possible to generate a wide variety of high-quality data that closely to actual anomalies.
-Liu et al.
-establish multistage GANs that decouple texture generation and background-anomaly fusion, effectively modeling both abnormal areas and contextual coherence.
-To address data scarcity,  Du et al.
-develop Con-GAN, featuring shared data augmentation and hypersphere-based loss, it simultaneously prevent overfitting and enhance anomaly diversity.
-Duan et al.
-also propose a two-stage GAN, namely DFMGAN, which consists of two processes: initial learning on normal samples and subsequent fine-tuning using abnormal samples.
-For complex anomaly synthesis,  Yang et al.
-innovatively combine large receptive fields for global structural modeling with patch-level refinement mechanisms, achieving synergistic integration of macro-context and micro-details.
+liu2019multistage establish multistage GANs that decouple texture generation and background-anomaly fusion, effectively modeling both abnormal areas and contextual coherence.
+To address data scarcity,  du2022new develop Con-GAN, featuring shared data augmentation and hypersphere-based loss, it simultaneously prevent overfitting and enhance anomaly diversity.
+duan2023few also propose a two-stage GAN, namely DFMGAN, which consists of two processes: initial learning on normal samples and subsequent fine-tuning using abnormal samples.
+For complex anomaly synthesis,  yang2025defect innovatively combine large receptive fields for global structural modeling with patch-level refinement mechanisms, achieving synergistic integration of macro-context and micro-details.
 Strengths and Weaknesses.
@@
 Additionally, as these models generate entire images rather than directly modifying normal samples, they struggle to preserve fine-grained structural details and contextual consistency.
-Full-image translation is a key approach within GM-Based Models, it synthesizes anomalies by transforming normal images via domain translation techniques like CycleGAN [Zhu et al.] and Pix2PixGAN [Isola et al.].
+Full-image translation is a key approach within GM-Based Models, it synthesizes anomalies by transforming normal images via domain translation techniques like CycleGAN [CycleGAN2017] and Pix2PixGAN [pix2pix].
 Full-image translation learns mappings between normal and anomaly domains, introducing targeted modifications (e.g.
@@
 Recent advancements significantly improve full-image translation by enhancing the fidelity and controllability of synthesized anomalies.
-Wen et al.
-extend CycleGAN within their anomaly detection pipeline, achieving superior performance.
-Niu et al.
-propose SDGAN, a CycleGAN-based framework incorporating additional discriminators to refine the translation process.
+wen2022new extend CycleGAN within their anomaly detection pipeline, achieving superior performance.
+niu2020defect propose SDGAN, a CycleGAN-based framework incorporating additional discriminators to refine the translation process.
 By focusing on modeling the distribution of abnormal samples, SDGAN enhances the realism of synthesized anomalies.
-Furthermore,  Zhang et al.
-introduce Defect-GAN, which explicitly models both the defacement and restoration processes rather than synthesizing anomalies arbitrarily.
+Furthermore,  zhang2021defect introduce Defect-GAN, which explicitly models both the defacement and restoration processes rather than synthesizing anomalies arbitrarily.
 Additionally, it leverages spatial distribution maps to preserve the appearance of normal backgrounds, ensuring structural consistency in generated images.
@@
 Recent advancements enhance local anomaly synthesis by aligning the mask with synthesized abnormal regions and improving spatial consistency between synthetic defects and their surrounding contexts.
-Zhang et al.
-enhance the synthesis of realistic anomalies by amplifying their divergence from normal patterns.
+zhang2024realnet enhance the synthesis of realistic anomalies by amplifying their divergence from normal patterns.
 Their proposed RealNet perturbs the variance of denoising diffusion during reverse diffusion to synthesize a global anomaly map.
-Wei et al.
-construct pseudo-normal backgrounds in abnormal regions to emphasize the distribution of anomalies.
-Wei et al.
-develop DCDGAN, which trains a model on anomaly-only textures and blends them with different backgrounds, achieving multi-class and diversified anomaly synthesis.
+wei2022mask construct pseudo-normal backgrounds in abnormal regions to emphasize the distribution of anomalies.
+wei2023diversified develop DCDGAN, which trains a model on anomaly-only textures and blends them with different backgrounds, achieving multi-class and diversified anomaly synthesis.
 Strengths and Weaknesses.
@@
 By capitalizing on abundant cues from various modalities, single-stage synthesis balances computational efficiency and domain specificity, making it an attractive solution for industrial applications.
-Sun et al.
-introduce a training-free anomaly synthesis method based on stable diffusion that maximizes alignment between abnormal regions and anomaly descriptions.
-Hu et al.
-generate embeddings from diverse modalities and adjust them dynamically during the diffusion process.
+sun2024cut introduce a training-free anomaly synthesis method based on stable diffusion that maximizes alignment between abnormal regions and anomaly descriptions.
+hu2024anomalyxfusion generate embeddings from diverse modalities and adjust them dynamically during the diffusion process.
 This strategy enables the creation of anomalies enriched with multimodal information, resulting in higher fidelity and diversity.
-Additionally,  Jiang et al.
-propose CAGEN, which employs mask annotations and text prompts to fine-tune ControlNet for precise control over the location and type of synthesized anomalies.
-Furthermore,  He et al.
-model cross-modal semantic features to allow fine-grained control over the characteristics of generated anomalies.
+Additionally,  jiang2024cagen propose CAGEN, which employs mask annotations and text prompts to fine-tune ControlNet for precise control over the location and type of synthesized anomalies.
+Furthermore,  he2024anomalycontrol model cross-modal semantic features to allow fine-grained control over the characteristics of generated anomalies.
 Strengths and Weaknesses.
@@
 Recent studies contribute notably to this field.
-Jin et al.
-propose a dual-branch VLM where one branch captures global context and the other targets abnormal regions, and it leverages a pre-trained segmentation for high-quality mask generation.
-Shi et al.
-integrate textual guidance into global information and local synthesis, introducing a novel mask algorithm that improves small anomaly synthesis and enables precise control over intensity and direction.
-Dai et al.
-introduce SeaS, a few-shot method that binds anomaly attributes to specific tokens using unbalanced anomaly prompts while aligning normal image features to preserve authenticity, it can also produce accurate mask annotations.
-Lastly,  Hu et al.
-develop AnomalyDiffusion, incorporating textual and positional information for precise anomaly localization and employing textual inversion to learn mask embeddings.
+jin2024dualanodiff propose a dual-branch VLM where one branch captures global context and the other targets abnormal regions, and it leverages a pre-trained segmentation for high-quality mask generation.
+shi2025few integrate textual guidance into global information and local synthesis, introducing a novel mask algorithm that improves small anomaly synthesis and enables precise control over intensity and direction.
+dai2024seas introduce SeaS, a few-shot method that binds anomaly attributes to specific tokens using unbalanced anomaly prompts while aligning normal image features to preserve authenticity, it can also produce accurate mask annotations.
+Lastly,  hu2024anomalydiffusion develop AnomalyDiffusion, incorporating textual and positional information for precise anomaly localization and employing textual inversion to learn mask embeddings.
 Strengths and Weaknesses.